Kog a brewer of a local beer called daa notes that when the price of daa reduces by GH5 the quantity demanded for daa increases by 5 gallons whereas the quantity supplied of daa reduces by 20 gallons. She indicated that the maximum quantity of daa demand and supply are 1500 gallons and negative 600 gallons respectively when the price of daa falls to zero.a) find the equilibrum price and quantity.
1
Expert's answer
2017-05-18T06:41:10-0400
When the price of daa reduces by GH5 the quantity demanded for daa increases by 5 gallons whereas the quantity supplied of daa reduces by 20 gallons. The maximum quantity of daa demanded and supplied are 1500 gallons and -600 gallons respectively when P = 0.
a) The price elasticity of demand is Ed = -5/5 = -1, the price elasticity of supply is Es = 20/5 = 4.
So, Pd = -Q + 1500, Ps = 4Q - 600.
In equilibrium Pd = Ps and Qd = Qs, so:
-Q + 1500 = 4Q - 600,
5Q = 2100,
The equilibrium quantity is Q = 420 units.
The equilibrium price is P = -420 + 1500 = GH1080.